OpenStudy (anonymous):

If f(x)=sigma (sin^2 (x))^k from k=1 to infinity, then f(1) is?

OpenStudy (anonymous):

@Hero @zepdrix @abb0t @Compassionate

OpenStudy (anonymous):

@tkhunny @beccaboo333 @Destinymasha

OpenStudy (beccaboo333):

I am not good with this kind of math. I'm sorry :/

OpenStudy (anonymous):

It's alright.

OpenStudy (tkhunny):

The ONLY x appearing in the definition is int he sine function. Substitute in there and see what you get. \(\sin(1) = 0.84147\) \(\sin^{2}(1) = 0.70807\) That's it. Now, use all you know about Geometric Series and add it up!

OpenStudy (anonymous):

\[ \sum _{k=1}^{\infty } \sin ^{2 k}(x)=\tan ^2(x) \] Do you know how to do it?

OpenStudy (anonymous):

Your answer is \[ \tan^2(1)=2.42552 \]

OpenStudy (anonymous):

But the answer should be 2.400.

OpenStudy (anonymous):

Yes to the nearest 10th

OpenStudy (anonymous):

This is a multiple choice problem with the answer choices a) 0.369, b) 0.585, c) 2.400, d) 2.426, and e) 3.426 and the answer is D.

OpenStudy (anonymous):

select 2.400

OpenStudy (anonymous):

But how to get there?

OpenStudy (anonymous):

\[ \sin^2(x) + \sin^4(x) +\sin^6(x) + \cdots=\\ \sin^2(x)(1 + \sin^2(x) +\sin^2(x) + \cdots=\\ \sin^2(x)\frac 1{1- \sin^2(x)}=\frac {\sin^2(x)}{\cos^2(x)}=\tan^2(x) \] It is a geometric series. We need to assume \[\Large \left|\sin^2(x) \right|<1 \]

OpenStudy (anonymous):

Do you understand now?

OpenStudy (anonymous):

I understand now. Thanks for helping me in this late night. I appreciated it.